# A Formalism for the Transport and Matching of Coupled Beams in Accelerators

### v1

@@ V Solutions for Beam Characterization and Diagnostics
While we can design and match the optics in coupled dynamics, it is important to consider beam characterization and diagnostics methods for coupled beams and lattices; for example, to distinguish between eigenmode emittances for fully coupled lattices or special configurations such as Derbenev’s Adapter. It is challenging to come up with a general diagnostics method for extracting eigenmode emittances from observables such as beam sizes. From our studies, we have identified the coupling strength, $u$, as an important parameter; see Section III.2. Through matching procedures, we have found that $u=1/2$, corresponding to fully coupled configurations, imposes certain constraints on coupled functions. These simplifications lead to possible methods of extracting the eigenmode emittances from beam observables. Eigenmode emittances can be calculated using the full knowledge of the beam matrix as
@@ V.1 Coupled Beam, Uncoupled Lattice ($u=1/2$)
Applying the simplifications that arise from $u=1/2$ case, for a downstream transfer matrix that is uncoupled: $\beta_{1x}=\beta_{2x}=\beta_{1}$, $\beta_{1y}=\beta_{2y}=\beta_{2}$. For simplicity, the measurement can be taken at the beam waist location where $\alpha_{1x}=\alpha_{2x}=\alpha_{1}=0$ and $\alpha_{1y}=\alpha_{2y}=\alpha_{2}=0$. The initial location measurement is taken when the beam’s transverse cross-section is aligned with $x-y$ plane with phases of coupling $\nu_{1,2}=\pi/2$. If the downstream lattice does not include any coupling elements, then the coupling strength parameter is always conserved, $u=1/2$. This leads to the important rms quantities:
@@ V.2 Extraction of Coupling Strength
The coupling strength stays constant if the transfer matrix is uncoupled. Depending on the coupling strength and periodicity of the phases of coupling, Fourier analysis applied to particle trajectories would reveal two peaks in the frequency diagram. For instance, for $u=1/2$, both modes contribute to the dynamics, which leads to two distinct peaks in the Fourier spectra, as illustrated in Fig. 14. The two peaks correspond to the tunes of modes $1$ and $2$, where the amplitudes of the peaks directly relate to the eigenmode emittances. The Fourier transformation of the phase space coordinate yields:
@@
Figure 14: FFT analysis of $u=1/2$ with $\nu_{1,2}=\pi/2$ beam in the $x$ plane. Equal eigenmode emittances (left), and factor of $100$ difference between eigenmode emittances (right). Tracked with ImpactX code in a simple ring design.
VI Conclusion
@@
In conclusion, we have presented an explicit formalism that governs the transformation of coupled beam optics functions in a similar way to the transformation of uncoupled beam Twiss functions. The coupled optics are derived from both Mais-Ripken parametrization and Lebedev-Bogacz eigenvector parametrization. We have shown the connection and equivalence between the Mais-Ripken and Lebedev-Bogacz parameters. Using the transfer matrix method with the generating vectors as the basis, it is straightforward to create an optimization routine to find the optimal periodic functions. Using a method for matching and optimizing coupled functions, we have shown that a fully coupled lattice can be created. The capability of producing matched coupled lattices will help explore new options for future accelerator and collider designs that could benefit from potential advantages in the unexplored area of coupled dynamics, such as the application of circular-mode beams and what they might offer. We have also presented a method for extracting eigenmode emittances from observables that could be measured in fully coupled lattices. Having a transfer matrix method for coupled optics functions enables the integration of this method into well-known design codes. It could also enable future AI/ML modeling for coupled lattice design, which would eventually lead to interesting new designs that we have not yet thought of.
VII Acknowledgments
@@ Appendix A Phase Advances From Coupled Transfer Matrix
$\begin{split}\vec{z}_{1}&=\begin{pmatrix}x_{1}\\ x^{\prime}_{1}\\ y_{1}\\ y^{\prime}_{1}\end{pmatrix}=\begin{pmatrix}\sqrt{\beta_{1x}}\cos\mu_{1}\\ \frac{-\alpha_{1x}\cos\mu_{1}-(1-u)\sin\mu_{1}}{\sqrt{\beta_{1x}}}\\ \sqrt{\beta_{1y}}\cos(\mu_{1}-\nu_{1})\\ \frac{-\alpha_{1y}\cos(\mu_{1}-\nu_{1})-u\sin(\mu_{1}-\nu_{1})}{\sqrt{\beta_{1y}}}\end{pmatrix},\qquad\vec{z}_{2}=\begin{pmatrix}x_{2}\\ x^{\prime}_{2}\\ y_{2}\\ y^{\prime}_{2}\end{pmatrix}=\begin{pmatrix}\sqrt{\beta_{1x}}\sin\mu_{x}\\ \frac{(1-u)\cos\mu_{1}-\alpha_{1x}\sin\mu_{1}}{\sqrt{\beta_{1x}}}\\ \sqrt{\beta_{1y}}\sin(\mu_{1}-\nu_{1})\\ \frac{u\cos(\mu_{1}-\nu_{1})-\alpha_{1y}\sin(\mu_{1}-\nu_{1})}{\sqrt{\beta_{1y}}}\end{pmatrix},\\ \vec{z}_{3}&=\begin{pmatrix}x_{3}\\ x^{\prime}_{3}\\ y_{3}\\ y^{\prime}_{3}\end{pmatrix}=\begin{pmatrix}\sqrt{\beta_{2x}}\cos(\mu_{2}-\nu_{2})\\ \frac{-\alpha_{2x}\cos(\mu_{2}-\nu_{2})-u\sin(\mu_{2}-\nu_{2})}{\sqrt{\beta_{2x}}}\\ \sqrt{\beta_{2y}}\cos\mu_{2}\\ \frac{-\alpha_{2y}\cos\mu_{2}-(1-u)\sin\mu_{2}}{\sqrt{\beta_{2y}}}\end{pmatrix},\qquad\vec{z}_{4}=\begin{pmatrix}x_{4}\\ x^{\prime}_{4}\\ y_{4}\\ y^{\prime}_{4}\end{pmatrix}=\begin{pmatrix}\sqrt{\beta_{2x}}\sin(\mu_{2}-\nu_{2})\\ \frac{u\cos(\mu_{2}-\nu_{2})-\alpha_{2x}\sin(\mu_{2}-\nu_{2})}{\sqrt{\beta_{2x}}}\\ \sqrt{\beta_{2y}}\sin\mu_{2}\\ \frac{(1-u)\cos\mu_{2}-\alpha_{2y}\sin\mu_{2}}{\sqrt{\beta_{2y}}}\end{pmatrix}.\end{split}$ (56)
@@ Appendix B Coupled Beam Matrix and Optics Functions
In order to have no tilt in $x-y$ plane, $\sigma_{xy}=0$. In coupled optics, no tilt suggests the phases of coupling are $\pi/2$.
@@ Appendix C Symmetric Alternating Gradient Design Condition
Symmetric alternating gradient (SAG) design is a midpoint plane symmetry that ensures equal phase advance transformation for quadrupole-based periodic cells. The design principle is shown in Fig. 15. The magnetic elements to the left of the middle of the cell are identical in strength but opposite in polarity, $K=-K$, with respect to the elements to the right. The transfer matrices are expressed as
$\begin{split}\mathcal{M}(0,L/2)=\begin{pmatrix}M&0\\ 0&N\end{pmatrix},\qquad\mathcal{M}(L,L/2)=\begin{pmatrix}N&0\\ 0&M\end{pmatrix}.\end{split}$ (63)

### v2

@@ V Solutions for Beam Characterization and Diagnostics
While we can design and match the optics in coupled dynamics, it is important to consider beam characterization and diagnostics methods for coupled beams and lattices; for example, to distinguish between eigenmode emittances for fully coupled lattices or special configurations such as Derbenev’s Adapter. It is challenging to come up with a general diagnostics method for extracting eigenmode emittances from observables such as beam sizes. From our studies, we have identified the coupling strength, $u$, as an important parameter; see Section III.2. Through matching procedures, we have found that $u=1/2$, corresponding to fully coupled configurations, imposes certain constraints on coupled functions. These simplifications lead to possible methods of extracting the eigenmode emittances from beam observables. Eigenmode emittances can be calculated generally using the full knowledge of the beam matrix as shown in Appendix B, which yields:
@@ V.1 Coupled Beam, Uncoupled Lattice ($u=1/2$)
Applying the simplifications that arise from $u=1/2$ case, for a downstream transfer matrix that is uncoupled: $\beta_{1x}=\beta_{2x}=\beta_{1}$, $\beta_{1y}=\beta_{2y}=\beta_{2}$. For simplicity, the measurement can be taken at the beam waist location where $\alpha_{1x}=\alpha_{2x}=\alpha_{1}=0$ and $\alpha_{1y}=\alpha_{2y}=\alpha_{2}=0$. The initial location measurement is taken when the beam’s transverse cross-section is aligned with xyx-y plane with phases of coupling $\nu_{1,2}=\pi/2$. If the downstream lattice does not include any coupling elements, then the coupling strength parameter is always conserved, $u=1/2$. This leads to the important rms quantities:
@@ V.2 Extraction of Coupling Strength
The coupling strength stays constant if the transfer matrix is uncoupled. The projections from mode $1$ and mode $2$ onto $x$ and $y$ planes will yield two Fourier peaks as a result of two modes of oscillations projecting onto real space as illustrated in Fig. 14. The two peaks correspond to the tunes of modes $1$ and $2$, where the relative amplitudes of the peaks reflect the eigenmodes emittances ratio. Although Fig. 14 was produced for $u=1/2$, the behavior is general. The Fourier transformation of the phase space coordinate yields:
@@
Figure 14: FFT analysis of $u=1/2$ with $\nu_{1,2}=\pi/2$ beam in the $x$ plane. Equal eigenmode emittances (left), and factor of 100100 difference between eigenmode emittances (right). Tracked with ImpactX code in a simple ring design.
VI Linear Difference Resonance
The use of skew quadrupoles adds two more resonances to the linear resonance conditions, namely sum and difference resonances. The difference resonance relates to the condition when the tunes satisfy: $Q_{x}-Q_{y}=n$ for $n$ an integer. The difference resonance is introduced through a resonance proximity parameter and skew quadrupole coefficient in the literature, $\Delta=Q_{x}-Q_{y}-k$. The averaged slow coordinate rotation vectors yield the equation:
$\frac{d}{dn}\begin{pmatrix}a\\ b\end{pmatrix}=-i\begin{pmatrix}\Delta&C\\ C&-\Delta\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}$, (56)
where $a,b$ are resonant frame vectors, $\Delta$ the resonance proximity parameter and $C$ the resonance driving term as shown in Appendix. D. The decoupling angle is computed with respect to the matrix shown above and yields:
$\tan 2\alpha=\frac{\Delta}{C}.$ (57)
Since the angle decouples the normalized motion, it can be associated with the Edwards-Teng angle defined in Section III.1. The decoupling angle can be related to the coupling strength parameter, $u$, with the definition $\tan^{2}\phi=u/(1-u)$, which yields the solution for the strength parameter as:
$u=\frac{1}{2}(1-\frac{\Delta}{\sqrt{\Delta^{2}-C^{2}}}).$ (58)
As we see, $\Delta=0$ corresponds to on-resonance behavior, which yields $u=1/2$ as the coupling strength. This is when two emittances get “shared” which also corresponds to the fully coupled scenario in the coupled parametrization. The shared emittance is seen from the projection emittances from Eq. (40), as for $u=1/2$ case reduces to:
$\epsilon_{x,y}=\frac{1}{2}(\epsilon_{1}+\epsilon_{2}),$ (59)
where $\epsilon_{1,2}$ are the eigenmode emittances intrinsic to the beam and $\epsilon_{x,y}$ the projected emittances. When $u=0$, the projected emittances are assigned the correct mode representations. The turn-by-turn emittance exchange rate is given in terms of resonance parameters in [18]. However, since the coupling strength parameter depends on the resonance parameters, the exchange rate can be expressed in terms of the coupling strength parameter of the distribution, since the projected emittances can be written as:
$\begin{split}\epsilon_{x}(n)&=\epsilon_{1}-u_{b}(n)(\epsilon_{1}-\epsilon_{2}),\\ \epsilon_{y}(n)&=\epsilon_{2}+u_{b}(n)(\epsilon_{1}-\epsilon_{2}),\end{split}$ (60)
where $u_{b}$ is the beams coupling strength and the rate of change of projected emittances with respect to turn number depends on the coupling strength parameter as:
$\begin{split}\frac{d\epsilon_{x}}{dn}&=-\frac{du_{b}}{dn}(\epsilon_{1}-\epsilon_{2}),\\ \frac{d\epsilon_{y}}{dn}&=\frac{du_{b}}{dn}(\epsilon_{1}-\epsilon_{2}).\end{split}$ (61)
As the beam continuously goes through the skew perturbation, the coupling strength parameter of the beam changes. Therefore, we can write an ansatz for the evolution of the coupling strength parameter as:
$u_{b}(n)=\frac{1}{2}(1-\cos(2\phi_{0}+2\pi n\chi)),$ (62)
where $\phi_{0}$ is initial phase relation where $\chi=\sqrt{\Delta^{2}+C^{2}}$ term. The propagation of the coupling strength parameter is then written as:
$u_{b}(n)=\frac{1}{2}+(u_{b0}-\frac{1}{2})\cos(2\pi n\chi)+\sqrt{u_{b0}(1-u_{b0})}\sin(2\pi n\chi)$ (63)
This results in the propagation of the projected emittances as:
$\begin{split}\epsilon_{x}(n)&=\frac{\epsilon_{1}}{2}+\frac{\epsilon_{2}}{2}-\left((u_{b0}-\frac{1}{2})\cos(2\pi n\chi)+\sqrt{u_{b0}(1-u_{b0})}\sin(2\pi n\chi)\right)(\epsilon_{1}-\epsilon_{2}),\\ \epsilon_{y}(n)&=\frac{\epsilon_{1}}{2}+\frac{\epsilon_{2}}{2}+\left((u_{b0}-\frac{1}{2})\cos(2\pi n\chi)+\sqrt{u_{b0}(1-u_{b0})}\sin(2\pi n\chi)\right)(\epsilon_{1}-\epsilon_{2}).\end{split}$ (64)
Where $\epsilon_{1,2}$ are invariant eigenmode emittances, $u_{b0}$ is initial coupling strength of the beam. Parametrizing the linear difference resonance from the perspective of the coupling strength parameter suggests that it is the coupling strength parameter of the beam that changes, rather than the emittances, which yields the exchange principle. An example of this dynamic behavior is observed in the generating function as it tracks through a linear periodic system with equal tunes $Q_{x}=Q_{y}$, featuring a skew perturbation at the system’s end. The results are shown in Fig. 15, where the coupled optics functions illustrate the behavior of the difference resonance. The off-mode beta functions start from zero, indicating no coupling, and become equal at the $u=1/2$ location. The on-mode functions tend to zero as the coupling strength parameter reaches one and periodically oscillate in that pattern. The minimum emittance exchange is fully realized when the coupling strength parameter $u$ reaches one, and the RMS emittance in $x$ swaps with the RMS emittance in $y$. As we see from the figure, the eigenmode emittances are invariant in this process due to the linearity of the transformation. The RMS projected emittances and eigenmode emittances are tracked and computed using ImpactX simulation [23].
Figure 15: Tracking through a periodic lattice with skew perturbation. Coupled beta functions(top-left), coupling strength parameter(top-right),RMS projected emittances(bottom-left), RMS eigenmode emittances(bottom-right)
VII Conclusion
@@
In conclusion, we have presented an explicit formalism that governs the transformation of coupled beam optics functions in a similar way to the transformation of uncoupled beam Twiss functions. The coupled optics are derived from both Mais-Ripken parametrization and Lebedev-Bogacz eigenvector parametrization. We have shown the connection and equivalence between the Mais-Ripken and Lebedev-Bogacz parameters. Using the transfer matrix method with the generating vectors as the basis, it is straightforward to create an optimization routine to find the optimal periodic functions. Using a method for matching and optimizing coupled functions, we have shown that a fully coupled lattice can be created. The capability of producing matched coupled lattices will help explore new options for future accelerator and collider designs that could benefit from potential advantages in the unexplored area of coupled dynamics, such as the application of circular-mode beams and what they might offer. We have also given a picture of linear difference resonance that arises from skew perturbation to a system and related the emittance exchange principle to the coupled optics functions with visualizations. We have also presented a method for extracting eigenmode emittances from observables that could be measured in fully coupled lattices. Having a transfer matrix method for coupled optics functions enables the integration of this method into well-known design codes. It could also enable future AI/ML modeling for coupled lattice design, which would eventually lead to interesting new designs that we have not yet thought of.
VIII Acknowledgments
@@ Appendix A Phase Advances From Coupled Transfer Matrix
[…]
@@ Appendix B Coupled Beam Matrix and Optics Functions
Using the normalization conditions of eigenvectors, $\vec{v}_{i}^{\dagger}S\vec{v}_{i}=-2i$, the strength of coupling can be deduced as:
$\begin{split}\kappa_{x}&=\sqrt{\frac{\beta_{2x}}{\beta_{1x}}},\quad\kappa_{y}=\sqrt{\frac{\beta_{1y}}{\beta_{2y}}},\\ A_{x}&=\alpha_{1x}\kappa_{x}-\alpha_{2x}\kappa_{x}^{-1},\\ A_{y}&=\alpha_{2y}\kappa_{y}-\alpha_{1y}\kappa_{y}^{-1},\\ u&=\frac{-\kappa_{x}^{2}\kappa_{y}^{2}\pm\sqrt{\kappa_{x}^{2}\kappa_{y}^{2}(1+\frac{A_{x}^{2}-A_{y}^{2}}{\kappa_{x}^{2}-\kappa_{y}^{2}}(1-\kappa_{x}^{2}\kappa_{y}^{2}))}}{1-\kappa_{x}^{2}\kappa_{y}^{2}}.\end{split}$ (69)
The normalization condition also allows one to extract single particle geometric amplitudes. Multiplying the phase space vector on the left hand side with $\vec{v}_{1}^{\dagger}S$ leads to deducing the single particle amplitude:
$\begin{split}2J_{1}&=\vec{v}_{1}^{\dagger}S\vec{z}\vec{z}^{T}S^{T}\vec{v}_{1}\\ &=(\frac{(1-u)^{2}+\alpha_{1x}^{2}}{\beta_{1x}})x^{2}+2\alpha_{1x}xx^{\prime}+\beta_{1x}x^{\prime 2}+(\frac{u^{2}+\alpha_{1y}^{2}}{\beta_{1y}})y^{2}+2\alpha_{1y}yy^{\prime}+\beta_{1y}y^{\prime 2}\\ &+\frac{2xy}{\sqrt{\beta_{1x}\beta_{1y}}}((u(1-u)+\alpha_{1x}\alpha_{1y})\cos\nu_{1}+(\alpha_{1y}-u(\alpha_{1x}+\alpha_{1y}))\sin\nu_{1})\\ &+2yx^{\prime}\sqrt{\frac{\beta_{1x}}{\beta_{1y}}}(\alpha_{1y}\cos\nu_{1}-u\sin\nu_{1})+2xy^{\prime}\sqrt{\frac{\beta_{1y}}{\beta_{1x}}}(\alpha_{1x}\cos\nu_{1}+(1-u)\sin\nu_{1})\\ &+2x^{\prime}y^{\prime}\sqrt{\beta_{1x}\beta_{1y}}\cos\nu_{1}.\end{split}$ (70)
The first two entries in the equation above show two Courant-Snyder like ellipses on $x-x^{\prime}$ and $y-y^{\prime}$ planes. The rest of the terms show the correlations between different phase planes that depend on the phases of coupling.
@@
In order to have no tilt in xyx-y plane, $\sigma_{xy}=0$. In coupled optics, no tilt suggests the phases of coupling are $\pi/2$.
@@
Under symplectic transformations the beam matrix transforms linearly, $\Sigma_{f}=\mathcal{M}^{T}\Sigma_{i}\mathcal{M}$. The 4D emittance is computed through the determinant of the beam matrix:
$det(\Sigma_{f})=det(\mathcal{M})det(\Sigma_{i})det(\mathcal{M})=\epsilon_{4D}^{2}=(\epsilon_{1}\epsilon_{2})^{2}.$ (74)
Since the product of emittances is known from the determinant of the beam matrix, a sum of emittances is needed so computing individual eigenmode emittances is possible. The diagonalized beam matrix will yield RMS eigenmode emittances as:
$\tilde{\Sigma}=T^{-1}\Sigma T=\mathrm{diag}(\epsilon_{1},\epsilon_{1},\epsilon_{2},\epsilon_{2}),$ (75)
where $T$ is a symplectic matrix that diagonalizes the beam matrix. Computing trace of the diagonalized beam matrix and symplectic unit matrix will yield the sum of RMS eigenmode emittances as:
$Tr[(\tilde{\Sigma}S)^{2}]=-2(\epsilon_{1}^{2}+\epsilon_{2}^{2}).$ (76)
Using the cyclic property of the trace and identities of the symplectic unit matrix and the diagonalizing matrix $T$, we can prove that the two traces are identical in value: $Tr[(\tilde{\Sigma}S)^{2}]=Tr[(\Sigma S)^{2}]=-2(\epsilon_{1}^{2}+\epsilon_{2}^{2})$. Knowing the sum of the two RMS eigenmode emittances and the product of them allows us to express each individual eigenmode emittances from the beam matrix, as for instance for mode 1:
[…]
$\begin{split}2\epsilon_{1}^{4}+\epsilon_{1}^{2}Tr[(\Sigma S)^{2}]+2det(\Sigma)=0,\end{split}$ (78)
the solution will yield:
$\epsilon_{1}=\frac{1}{2}\sqrt{-Tr[(\Sigma S)^{2}]+\sqrt{Tr^{2}[(\Sigma S)^{2}]-16det(\Sigma)}}.$ (79)
@@ Appendix C Symmetric Alternating Gradient Design Condition
Symmetric alternating gradient (SAG) design is a midpoint plane symmetry that ensures equal phase advance transformation for quadrupole-based periodic cells. The design principle is shown in Fig. 16. The magnetic elements to the left of the middle of the cell are identical in strength but opposite in polarity, K=KK=-K, with respect to the elements to the right. The transfer matrices are expressed as
[…]
@@
Appendix D Linear Difference Resonance Driven By Skew Quadrupole Perturbation
The linear difference resonance is observed when uncoupled Hamiltonian is perturbed by skew quadrupole potential, $V_{skew}=gxy$. The perturbed Hamiltonian is written as:
$H=Q_{x}J_{x}+Q_{y}J_{y}+V_{skew}(\psi_{x},\psi_{y},J_{x},J_{y}).$ (81)
Where, for the skew potential Courant-Snyder coordinate parametrization is used:
$H_{skew}=g\sqrt{J_{x}J_{y}}\cos(\Psi_{x}-\Psi_{y})+g\sqrt{J_{x}J_{y}}\cos(\Psi_{x}+\Psi_{y}),$ (82)
which yields the two conditions for sum and difference resonances. Introducing a change of variable, $G=\frac{1}{2}(J_{x}+J_{y})$ and $I=\frac{1}{2}(J_{x}-J_{y})$ transforms the Hamiltonian:
$H=\delta Q_{x}+\Delta I+\tilde{g}\sqrt{G^{2}-I^{2}}\cos\xi,$ (83)
where $\delta=Q_{x}+Q_{y}$ and $\Delta=Q_{x}-Q_{y}-m$. The angle $\xi$ depends on the difference of betatron phases, since there is no dependence on the sum of betatron phases, the Hamiltonians equations yield an invariant, $\dot{G}=0$, which is the sum of actions of $x$ and $y$ planes, the rate of change of actions depends on $\dot{I}$ which gives opposite signs: $\dot{J}_{x}=-\dot{J}_{y}$ which yields in the emittance exchange formalism. The equations of motion yields the Hill’s equations. The solution for the skew potential is carried out in normal form coordinates:
$h_{x}=\frac{1}{\sqrt{2}}(X+iP_{x})=\sqrt{J_{x}}e^{-i\Psi_{x}},\qquad h_{y}=\frac{1}{\sqrt{2}}(Y+iP_{y})=\sqrt{J_{y}}e^{-i\Psi_{y}}.$ (84)
The skew perturbation is written as: $\Delta p_{x}=gy$ and $\Delta p_{y}=gx$. Expressing the coordinates in terms of normalized coordinates and expanding to normal-form coordinates yields:
$\begin{split}\Delta h_{x}&=\frac{i}{2}g\sqrt{\beta_{x}\beta_{y}}(h_{x}+h_{x}^{*}),\\ \Delta h_{y}&=\frac{i}{2}g\sqrt{\beta_{x}\beta_{y}}(h_{y}+h_{y}^{*}).\end{split}$ (85)
One turn map rotates the normalized vectors, hence; the normal-form vectors are written as: $h_{x}(n)=h_{x}e^{-i2\pi nQ_{x}}$. Moving to resonant co-rotating frame by introducing slow variables:
$a_{n}=h_{x}e^{i2\pi Q_{x}n},\quad b_{n}=h_{y}^{*}e^{-i2\pi(Q_{y}+k)n},$ (86)
with detuning resonance proximity parameter as: $\Delta=Q_{x}-Q_{y}-k$, the averaging yields:
$\frac{d}{dn}\begin{pmatrix}a\\ b\end{pmatrix}=-i\begin{pmatrix}\Delta&C\\ C&-\Delta\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}$ (87)
Which yields the decoupling angle as:
$\tan 2\alpha=\frac{\Delta}{C},$ (88)
Where, $C$ is defined as the driving term:
$C=\frac{1}{2\pi}\oint g(s)\sqrt{\beta_{x}\beta_{y}}e^{i(\Psi_{x}-\Psi_{y}-k\theta)},$ (89)
for $\theta=2\pi s/C$.